OpenStudy (darthvader2900):

Solve for x. x+2/3 = -4 A. 1 B. –3 C. –10 D. –14

OpenStudy (igreen):

Subtract 2/3 to both sides, what's -4 - 2/3?

OpenStudy (darthvader2900):

@iGreen I think the answer is B.

OpenStudy (igreen):

Oh

OpenStudy (igreen):

Wait, it's like this? \(\sf \dfrac{x + 2}{3} = -4\)

OpenStudy (darthvader2900):

Yeah

OpenStudy (igreen):

Okay, first step is to multiply 3 to both sides, -4 * 3 = ?

OpenStudy (darthvader2900):

-12

OpenStudy (darthvader2900):

I believe that the answer is B.-3

OpenStudy (darthvader2900):

hello?

OpenStudy (igreen):

We subtract 2 to both sides now. \(\sf x + 2 = -12\) -12 - 2 = ?

OpenStudy (darthvader2900):

-14

OpenStudy (darthvader2900):

ohhh the answer is D. thanx

OpenStudy (igreen):

Yep